# A nonequilibrium quantum Otto engine enhanced via multi-parameter control

### v1

@@ Abstract
Advances in experimental control of interacting quantum many-body systems with multiple tunable parameters—such as ultracold atomic gases and trapped ions—are driving rapid progress in quantum thermodynamics and enabling the design of quantum thermal machines. In this work, we utilize a sudden quench approximation as a means to investigate the operation of a quantum thermodynamic Otto cycle in which multiple parameters are simultaneously controllable. The method applies universally to many-body systems where such control is available, and therefore provides general principles for investigating their operation as a working medium in quantum thermal machines. We investigate application of this multi-parameter quench protocol in an experimentally realistic one-dimensional Bose gas as the working fluid, with control over both the frequency of an external harmonic trap and the interparticle interaction strength. We derive a general inequality for the net work of this two-parameter Otto cycle, demonstrating that this protocol out-performs its constituent single-parameter Otto cycles when operating as an engine, and additionally implying an enhancement to the coefficient of performance when operating as a refrigerator. Further, we demonstrate that multi-parameter control can exhibit dramatically improved performance of the Otto engine when compared not only to single-parameter constituent quenches but also to the combined effect of its constituent engine cycles.
@@ I Introduction
Engine cycles have been central to the historical development of quantum thermodynamics, with origins dating back to 1953 in Scovil and Schulz-Dubois’ analysis of the maser as a single-body 3-level quantum heat engine [15]. Recently, in order to advance the understanding of quantum engines and their scaling, many researchers have focused on the operation of quantum devices in quantum many-body systems [16, 17, 6, 18, 19, 20, 21, 22, 23, 24, 25, 26, 27, 28, 29]. Such many-body interacting quantum systems enable the investigation of the role of uniquely quantum effects in engine operation, such as quantum coherence [30, 31, 32, 33, 34, 35] or correlations [36, 37, 38, 39, 19].
Notable recent examples have been the experimental realizations of quantum Otto engines in interacting ultracold Bose gases [16, 17]. These experiments exploited quasistatic control over both interparticle interaction strength and the frequency of an external harmonic trap in an alternating fashion, thus realizing uniquely quantum many-body thermodynamic devices. In particular, Ref. [16] elegantly demonstrated the role that quantum statistics played in the performance of their quasistatic quantum Otto engine cycle. Extending the investigation of quantum many-body engine cycles to non-quasistatic protocols, i.e. to out-of-equilibrium regimes of operation, is essential in order to understand their performance in scenarios that generate a finite power output, which vanishes in the quasistatic limit. However, such an extension remains challenging due to the typical complexity of simulating the out-of-equilibrium dynamics of interacting many-body systems in experimentally realistic parameter regimes, where precise and rapid control is available over various system parameters [40, 41, 42, 43, 28, 44, 20]. An extreme version of an out-of-equilibrium engine operation is realized via a sudden quench, where one can approximate the final post-quench state as unchanged from its initial thermal equilibrium state in order to again rely on exact thermal equilibrium expectation values for calculation of the net work and efficiency [19, 21].
In this work, we examine a quantum Otto engine cycle, as well as other related thermal machines, such as an Otto refrigerator, operating under a sudden quench of multiple externally tunable parameters. In particular, we extend the recent work on Otto engine cycles, operating under a single parameter quench for arbitrary quantum models in Ref. [21], to the case of simultaneous quenching of multiple control parameters. We apply this formalism to a harmonically trapped one-dimensional (1D) Bose gas, where we explore a simultaneous sudden quench of the harmonic trapping frequency and the interatomic interaction strength. This experimentally realizable engine cycle is shown to possess a region of dramatically enhanced performance compared not only to when quenching only a single parameter, but also to the combined performance under two single-parameter quenches.
@@ II Multi-parameter sudden quench Otto engine
We begin by describing a general physical model that may correspond to, e.g., an ultracold quantum gas or a spin chain on a lattice, which incorporates multiple externally controllable parameters. In detail, in the context of second quantized quantum mechanics, we define a model Hamiltonian that consists first of the operators $\{\hat{\mathcal{V}}^{(\alpha)}\}$ (indexed by $\alpha$) intended to correspond to all operators with external control over a scalar strength parameter, denoted $\{c^{(\alpha)}\}$, such that the Hamiltonian contains the terms $c^{(\alpha)}\hat{\mathcal{V}}^{(\alpha)}$. The second set of operators, collectively denoted $\hat{H}_{0}$, corresponds to any and all terms without externally controlled strength parameters, a common example of which is the kinetic energy. Thus, the full Hamiltonian is given by,
@@
where we used the conservation of energy W+E=0W\!+\!E\!=\!0, with $E=E_{\mathbf{B}\to\mathbf{C}}\!+\!E_{\mathbf{D}\to\mathbf{A}}$ being the total energy. Importantly, the energy exchanged with the reservoirs during the equilibration strokes may take any form (e.g. heat, chemical work, etc.), requiring only that the set of strength parameters, $\{c^{(\alpha)}\}$, are kept constant. The case where energy intake and output is given entirely by heat corresponds to a conventional quantum Otto heat engine cycle, in which case the generalized efficiency is equivalent to the standard thermodynamic efficiency [45, 46]. Non-standard, or generalized, efficiency is utilized in cycles where the system-reservoir contact contains non-thermal components. A simple example of non-thermal contact is diffusive contact, which may be achieved through an exchange of particles between the system and reservoirs during the equilibration strokes; such a device is typically referred to as a thermo-chemical Otto engine [47, 48, 20, 21].
@@
In the following we explore a particular case of this multi-parameter sudden quench Otto cycle in an experimentally realizable harmonically trapped 1D Bose gas. Such a system is amenable to rapid experimental control over both the frequency of the external harmonic trap and over the strength of interparticle interactions [7, 49, 50, 51, 26]. However, we emphasize here that the formalism outlined in this section applies universally to other quantum models capable of rapid control over multiple external parameters [21].
III The 1D Bose gas
@@
The Lieb-Liniger model of a 1D Bose gas is a paradigmatic quantum model describing an ultracold atomic gas with contact pairwise interactions, being both experimentally realizable and having exact theoretical results [52, 53, 54, 7, 55, 56, 49, 54, 57, 58]. This model, in the uniform limit, consists first of an uncontrolled Hamiltonian operator, $\hat{H}_{0}$, corresponding to the kinetic energy,
@@
$c^{(2)}\hat{\mathcal{V}}^{(2)}=c^{(2)}\hat{\overline{G_{2}}}=c^{(2)}\!\int dx\hat{G}_{2}(x),$ (7)
where the two-body operator $\hat{\mathcal{V}}^{(2)}\equiv\hat{\overline{G_{2}}}=\int dx\hat{G}_{2}(x)$ corresponds to the integrated local (same-point) pair correlation $\hat{G}_{2}(x)\equiv\hat{G}_{2}(x,x)\!=\!\hat{\Psi}^{\dagger}(x)\hat{\Psi}^{\dagger}(x)\hat{\Psi}(x)\hat{\Psi}(x)$, which we note does not depend on $x$ in a uniform system due to the translational invariance. The interaction strength $c^{(2)}$ is given by $c^{(2)}\!\equiv\!g/2\simeq\!\hbar\omega_{\perp}a_{s}$ away from confinement induced resonances [59] and is experimentally controllable via the tight transverse trapping frequency, $\omega_{\perp}$, or magnetic Feshbach resonance [60] tuning of the 3D $s$-wave scattering length $a_{s}$. We may utilize the normalized local (same-point) two-body correlation function,
@@
$c^{(1)}\hat{\mathcal{V}}^{(1)}\equiv\hat{V}=\frac{1}{2}m\omega^{2}\int dx\,x^{2}\hat{\rho}(x),$ (9)
with $\hat{\mathcal{V}}^{(1)}\!\equiv\!\frac{1}{2}m\int dx\,x^{2}\hat{\rho}(x)$, where $c^{(1)}\!\equiv\!\omega^{2}$ corresponds to the harmonic trapping frequency squared and $\hat{\rho}(x)\!=\!\hat{\Psi}^{\dagger}(x)\hat{\Psi}(x)$ is the particle number density operator. We additionally define the atomic position variance, $\langle x^{2}\rangle\!\equiv\!\int dx\,x^{2}\langle\hat{\rho}(x)\rangle$, such that $\langle\hat{\mathcal{V}}^{(1)}\rangle\!=\!m\omega^{2}\langle x^{2}\rangle/2$.
In the presence of an external trapping potential, one may again utilize the numerically exact TBA introduced above, along with the additional assumption of a local density approximation [66], to accurately model the non-uniform system at finite temperature. Through the external harmonic trap, the density profile gains a dependence on the position, $\rho\!=\!\rho(x)$. The gas may then be described in terms of the dimensionless interaction strength and temperature at the trap center, i.e. $\gamma_{0}\!=\!mg/\hbar^{2}\rho(0)$ and $\tau_{0}\!=\!2mk_{B}T/\hbar^{2}\rho(0)^{2}$, and the average total number of particles in the system, $N\!=\!\int dx\langle\hat{\rho}(x)\rangle$ [66]. The interaction-driven and separately the volumetric (i.e. harmonic frequency quench) quantum Otto cycles for this system have been previously investigated under both quasi-static and sudden quench protocols [20, 21, 27, 17, 16].
@@
IV Two-parameter enhancement
Before turning to calculation of the net work, we examine the properties of this two-parameter Otto cycle in comparison with the two single-parameter Otto cycles, taken separately. These single-parameter quantum Otto cycles correspond to those commonly considered in thermodynamics [21], consisting of control over only the ratio $g_{h}/g_{l}$ (or $\omega_{h}^{2}/\omega_{l}^{2}$) while keeping the complementary quantity constant, i.e., $\omega_{h}^{2}/\omega_{l}^{2}\!=\!1$ (or $g_{h}/g_{l}\!=\!1$). Through this, we will demonstrate that the net work of the two-parameter Otto cycle is always greater than the sum of the single-parameter cycles taken in isolation. We then generalize the arguments presented for the harmonically trapped 1D Bose gas to general quantum models describing scenarios where the two-parameter cycle out-performs the sum of the single-parameter in terms of net work.
For clarity, we emphasize the dependence of the equilibrium states $(h)$ and $(l)$ on the interaction strength and harmonic trapping frequency as $\langle\cdot\rangle_{(g,\omega)}$. To illustrate this notation, we show the total system energy, $\langle\hat{H}\rangle_{(g,\omega)}$, in Fig. 1, and additionally demonstrate the path in parameter space which both the single-parameter and two-parameter Otto cycles traverse. Using this notation, the work input ($W_{\mathbf{A}\to\mathbf{B}}$) of the interaction-driven single-parameter Otto cycle is
$W_{\mathbf{A}\to\mathbf{B}}^{g}=$ $\frac{1}{2}(g_{h}-g_{l})\langle\hat{\overline{G_{2}}}\rangle_{(g_{l},\omega_{l})}.$ (11)
Similarly, the work input of the volumetric single-parameter Otto cycle is
$W_{\mathbf{A}\to\mathbf{B}}^{\omega}=$ $\frac{1}{2}m(\omega_{h}^{2}-\omega_{l}^{2})\langle x^{2}\rangle_{(g_{l},\omega_{l})}.$ (12)
Note that we have chosen to begin all cycles in the same low energy equilibrium state, characterized by the same temperature, particle number, and set of parameters $g_{l}$ and $\omega_{l}$.
Figure 1: Total system energy ($\langle\hat{H}\rangle_{(g,\omega)}$) diagram for the sudden quench Otto cycle with a harmonically trapped 1D Bose gas as the working fluid. Here, control is over both the interaction strength, $g$, and the external harmonic trapping frequency, $\omega$, given on the horizontal and vertical axes, respectively, for realistic experimental values [68]. Beginning in the low energy equilibrium state, given by $(g_{l},\omega_{l})$, one may consider a single-parameter Otto cycle where either interaction strength is suddenly quenched, $(g_{l},\omega_{l})\!\to\!(g_{h},\omega_{l})$, corresponding to an interaction-driven Otto cycle, or the harmonic trapping frequency is quenched, $(g_{l},\omega_{l})\!\to\!(g_{l},\omega_{h})$, corresponding to a volumetric Otto cycle. In contrast, the two-parameter Otto cycle consists of quenching both parameters, $(g_{l},\omega_{l})\!\to\!(g_{h},\omega_{h})$, resulting in an enhanced performance (see text) over the sum of the single-parameter Otto cycles.
We are interested in the difference in net work between the two-parameter Otto cycles and the sum of the single-parameter Otto cycles, which we denote
$\Delta W=(-W)-(-W^{g}-W^{\omega}).$ (13)
For the case of engine operation, having $\Delta W\!>\!0$ implies that the net work generated by two-parameter Otto cycle is greater than the sum of the single-parameter cycles taken together.
Inspecting first the work input from the compression stroke, $W_{\mathbf{A}\to\mathbf{B}}$, of the two-parameter Otto cycle,
$W_{\mathbf{A}\to\mathbf{B}}\simeq$ $\frac{1}{2}(g_{h}-g_{l})\langle\hat{\overline{G_{2}}}\rangle_{(g_{l},\omega_{l})}$
$+\frac{1}{2}m(\omega_{h}^{2}-\omega_{l}^{2})\langle x^{2}\rangle_{(g_{l},\omega_{l})},$ (14)
we observe that this is exactly the sum of the work inputs of the interaction-driven and volumetric single-parameter Otto cycles, given in Eqs. (11) and (12), respectively. This results from the fact that, regardless of cycle type, the system begins in the same low energy equilibrium state and implies that $\Delta W>0$ cannot originate from the compression stroke.
Therefore, a nonzero difference in net work $\Delta W$, if any, can only originate from the difference in the work outputs in the expansion stroke, $W_{\mathbf{C}\to\mathbf{D}}$, of these Otto engine cycles,
$-\Delta W=(-W_{\mathbf{C}\to\mathbf{D}})-(-W_{\mathbf{C}\to\mathbf{D}}^{g}-W_{\mathbf{C}\to\mathbf{D}}^{\omega}).$ (15)
For the two-parameter Otto cycle, the work output in the expansion stroke is given by
$-W_{\mathbf{C}\to\mathbf{D}}$ $=\frac{1}{2}(g_{h}-g_{l})\langle\hat{\overline{G_{2}}}\rangle_{(g_{h},\omega_{h})}$
$\quad+\frac{1}{2}m(\omega_{h}^{2}-\omega_{l}^{2})\langle x^{2}\rangle_{(g_{h},\omega_{h})}.$ (16)
In contrast, work outputs of the single-parameter Otto cycles, which keep the complementary parameters constant, are given by
$W_{\mathbf{C}\to\mathbf{D}}^{g}=\frac{1}{2}(g_{h}-g_{l})\langle\hat{\overline{G_{2}}}\rangle_{(g_{h},\omega_{l})},$ (17)
for the interaction-driven single-parameter Otto cycle, and by
$W_{\mathbf{C}\to\mathbf{D}}^{\omega}=\frac{1}{2}m(\omega_{h}^{2}-\omega_{l}^{2})\langle x^{2}\rangle_{(g_{l},\omega_{h})}.$ (18)
for the volumetric single-parameter Otto cycle. Substituting Eqs. (IV), (17), and (18) into Eq. (15) for $\Delta W$, gives:

### v2

# Enhanced performance of sudden-quench quantum Otto cycles via multi-parameter control

@@ Abstract
Advances in experimental control of interacting quantum many-body systems with multiple tunable parameters—such as ultracold atomic gases and trapped ions—are driving rapid progress in quantum thermodynamics and enabling the design of quantum thermal machines. In this work, we utilize a sudden quench approximation as a means to investigate the operation of a quantum thermodynamic Otto cycle in which multiple parameters are simultaneously controllable. The method applies universally to many-body systems where such control is available, and therefore provides general principles for investigating their operation as a working medium in quantum thermal machines. We investigate application of this multi-parameter quench protocol in an experimentally realistic one-dimensional Bose gas, as well as in the transverse-field Ising model. We find that such a multi-parameter Otto cycle, when operating as an engine, outperforms not only its constituent single-parameter Otto cycles in terms of the net work and efficiency, but also the combined net work of its constituent engine cycles when added together independently. We also find that a similar multi-parameter enhancement applies to the coefficient of performance when the Otto cycle operates as a refrigerator.
@@ I Introduction
Engine cycles have been central to the historical development of quantum thermodynamics, with origins dating back to 1953 in Scovil and Schulz-Dubois’ analysis of the maser as a single-body 3-level quantum heat engine [15]. Recently, in order to advance the understanding of quantum engines and their scaling, many researchers have focused on the operation of quantum devices in quantum many-body systems [16, 17, 6, 18, 19, 20, 21, 22, 23, 24, 25, 26, 27, 28, 29]. Such many-body interacting quantum systems enable investigation into the role of uniquely quantum effects in engine operation, such as quantum coherence [30, 31, 32, 33, 34, 35] or correlations [36, 37, 38, 39, 19].
Notable recent examples have been the experimental realizations of quantum Otto engines in interacting ultracold Bose gases [16, 17]. These experiments exploited quasistatic control over both interparticle interaction strength and the frequency of an external harmonic trap in an alternating fashion, thus realizing uniquely quantum many-body thermodynamic devices. In particular, Ref. [16] elegantly demonstrated the role that quantum statistics played in the performance of their quasistatic quantum Otto engine cycle.
Extending the investigation of quantum many-body engine cycles to non-quasistatic protocols, i.e. to out-of-equilibrium regimes of operation, is essential in order to understand their performance in scenarios that generate a finite power output, which vanishes in the quasistatic limit. However, such an extension remains challenging due to the typical complexity of simulating the out-of-equilibrium dynamics of interacting many-body systems in experimentally realistic parameter regimes, where precise and rapid control is available over various system parameters [40, 41, 42, 43, 28, 44, 20]. An extreme version of an out-of-equilibrium engine operation is realized via a sudden quench, where one can approximate the final post-quench state as unchanged from its initial thermal equilibrium state in order to again rely on exact thermal equilibrium expectation values for calculation of the net work and efficiency [19, 21].
In this work, we examine a quantum Otto engine cycle, as well as other related thermal machines, such as an Otto refrigerator, operating under a sudden quench of multiple externally tunable parameters. In particular, we extend the recent work on Otto engine cycles, operating under a single parameter quench for arbitrary quantum models in Ref. [21], to the case of simultaneous quenching of multiple control parameters. We apply this formalism to the harmonically trapped one-dimensional (1D) Bose gas and the transverse-field Ising model (TFIM), where we observe a general enhancement to the net work under two-parameter control when contrasted with the combined performance of two single-parameter quenches.
@@ II Multi-parameter sudden quench Otto engine
We begin by describing a general physical model that may correspond to, e.g., an ultracold quantum gas or a spin chain on a lattice, which incorporates multiple externally controllable parameters. In detail, in the context of second quantized quantum mechanics, we define a model Hamiltonian that consists first of the operators $\{\hat{\mathcal{V}}^{(\alpha)}\}$ (indexed by $\alpha=1,2,3,$...) intended to correspond to all operators with external control over a scalar strength parameter, denoted $\{c^{(\alpha)}\}$, such that the Hamiltonian contains the terms $c^{(\alpha)}\hat{\mathcal{V}}^{(\alpha)}$. The second set of operators, collectively denoted $\hat{H}_{0}$, corresponds to any and all terms without externally controlled strength parameters, a common example of which is the kinetic energy. Thus, the full Hamiltonian is given by,
@@
where we used the conservation of energy W+E=0W\!+\!E\!=\!0, with $E=E_{\mathbf{B}\to\mathbf{C}}\!+\!E_{\mathbf{D}\to\mathbf{A}}$ being the total energy.
@@
Importantly, the energy exchanged with the reservoirs during the equilibration strokes may take any form (e.g. heat, chemical work, etc.), requiring only that the set of strength parameters, $\{c^{(\alpha)}\}$, are kept constant. The case where energy intake and output is given entirely by heat corresponds to a conventional quantum Otto heat engine cycle, in which case the generalized efficiency is equivalent to the standard thermodynamic efficiency [45, 46]. Non-standard, or generalized, efficiency is utilized in cycles where the system-reservoir contact contains non-thermal components. A simple example of non-thermal contact is diffusive contact, which may be achieved through an exchange of particles between the system and reservoirs during the equilibration strokes; such a device is typically referred to as a thermo-chemical Otto engine [47, 48, 20, 21].
@@
In the following we investigate the conditions under which this multi-parameter quantum Otto cycle exhibits an unambiguous benefit over single-parameter operation. Following this, in Secs. IV-VI, we explore a particular case of this multi-parameter sudden quench Otto cycle in an experimentally realizable harmonically trapped 1D Bose gas. Such a system is amenable to rapid experimental control over both the frequency of the external harmonic trap and over the strength of interparticle interactions [7, 49, 50, 51, 26]. Finally, in Sec. VII, we demonstrate application of multi-parameter Otto engine to the TFIM, additionally utilizing small parameter quenches, through which we explore the role of criticality for this two-parameter Otto cycle.
III Multi-parameter enhancement
For quantum systems where multiple strength parameters, $c^{(\alpha)}$, may be simultaneously suddenly quenched, one may generally consider the scenario where each parameter drives its own Otto cycle, while leaving all other parameters constant. From this perspective, it is interesting to consider whether operating the multi-parameter Otto cycle produces a benefit over the sum of these single-parameter cycles.
In order to examine the differences between the general multi-parameter Otto cycle and its single-parameter counterparts, we begin by defining the dependence of the equilibrium operator expectation values, $\langle\hat{\mathcal{V}}^{(\alpha)}\rangle$, on the strength parameters, $c^{(\alpha)}$. In particular, for a multi-parameter sudden quench Otto cycle where $M$ parameters are simultaneously quenched, it is useful to write the equilibrium expectation value of each operator explicitly as a function of all parameters,
$\langle\hat{\mathcal{V}}^{(\alpha)}\rangle\left(c^{(1)},c^{(2)},\dots,c^{(M)};T\right).$ (6)
Doing so highlights the fact that each expectation values depend on all parameters defining the equilibrium state, in addition to the temperature, $T$.
We may then define the expectation value of each operator within the Hamiltonian in the low energy equilibrium state,
$\langle\hat{\mathcal{V}}^{(\alpha)}\rangle_{l}\equiv\langle\hat{\mathcal{V}}^{(\alpha)}\rangle\left(c^{(1)}_{l},\dots,c^{(M)}_{l};T_{l}\right),$ (7)
where all controllable parameters and the temperature are taken at their low energy values, $c^{(\alpha)}_{l}$ and $T_{l}$. The high energy equilibrium state, $h$, may be described similarly,
$\langle\hat{\mathcal{V}}^{(\alpha)}\rangle_{h}\equiv\langle\hat{\mathcal{V}}^{(\alpha)}\rangle\left(c^{(1)}_{h},\dots,c^{(M)}_{h};T_{h}\right).$ (8)
Using these notations, we express the net work of the sub-cycle associated only with the single $c^{(\alpha)}\,\hat{\mathcal{V}}^{(\alpha)}$ term in the Hamiltonian,
$W^{(\alpha)}_{\mathrm{multi}}=\left(c^{(\alpha)}_{h}-c^{(\alpha)}_{l}\right)\left(\langle\hat{\mathcal{V}}^{(\alpha)}\rangle_{h}-\langle\hat{\mathcal{V}}^{(\alpha)}\rangle_{l}\right).$ (9)
The total net work of the multi-parameter quench, with all $c^{(\alpha)}$’s being quenched between their respective $l$ to $h$ values, is additive, and hence
$W_{\mathrm{multi}}=\sum_{\alpha=1}^{M}W^{(\alpha)}_{\mathrm{multi}}$. (10)
In contrast, one may define the single-parameter Otto cycle, generated via control over only a single strength parameter, $c^{(\alpha)}$, while leaving all others constant at their low energy values. An example of such an Otto cycle is the volumetric Otto cycle in an interacting 1D Bose gas, where only the frequency of the external harmonic trap is controlled, and the inter-particle interaction strength left constant. Otto cycles of this type were previously investigated in Ref. [21] for arbitrary quantum models under a sudden quench approximation.
In order to express the net work of the single-parameter Otto cycle associated with a quench of the control parameter $c^{(\alpha)}$ from $c^{(\alpha)}_{l}$ to $c^{(\alpha)}_{h}$, we define the high-energy equilibrium expectation value for the corresponding operator,
$\langle\hat{\mathcal{V}}^{(\alpha)}\rangle_{h(\alpha)}\!\equiv\!\langle\hat{\mathcal{V}}^{(\alpha)}\rangle\left(c^{(1)}_{l},\dots,c^{(\alpha)}_{h},\dots,c^{(M)}_{l};T_{h}\right).$ (11)
Importantly, all parameters except $c^{(\alpha)}$ are kept constant at their low energy ($l$) values. The fact that only $c^{(\alpha)}$ is being quenched is reflected by the use of subscript $h(\alpha)$ in $\langle\hat{\mathcal{V}}^{(\alpha)}\rangle_{h(\alpha)}$, instead of simply $h$ as in Eq. (8). The net work of this single-parameter quench Otto cycle, denoted by $W_{\mathrm{single}}^{(\alpha)}$, is then given by
$W^{(\alpha)}_{\mathrm{single}}=\left(c^{(\alpha)}_{h}-c^{(\alpha)}_{l}\right)\left(\langle\hat{\mathcal{V}}^{(\alpha)}\rangle_{h(\alpha)}-\langle\hat{\mathcal{V}}^{(\alpha)}\rangle_{l}\right).$ (12)
Compared to Eq. (9), the superscript $(\alpha)$ here refers to the quenched parameter $c^{(\alpha)}$ (from the $l$ to $h$ value, while all the others remain at their $l$ values), and the net work comes again from a single $c^{(\alpha)}\,\hat{\mathcal{V}}^{(\alpha)}$ term in the Hamiltonian. We define the total net work due to quenching all parameters $c^{(\alpha)}$ individually as $W_{\mathrm{single}}$,
$W_{\mathrm{single}}=\sum_{\alpha=1}^{M}W^{(\alpha)}_{\mathrm{single}}$. (13)
We may now compare the net work of the multi-parameter Otto cycle with the sum of the net work achieved by the single-parameter cycles by their difference,
$-\Delta W\equiv-W_{\mathrm{multi}}-\left(-W_{\mathrm{single}}\right),$ (14)
We define a multi-parameter enhancement to the net work by the condition $\Delta W\!>\!0$. We then substitute the expressions for the net work of both the multi- and single-parameter Otto cycles, given in Eq. (9) and Eq. (12), respectively, to find
$-\Delta W\!=\!\sum_{\alpha=1}^{M}\left(c^{(\alpha)}_{h}\!-\!c^{(\alpha)}_{l}\right)\left(\big\langle\hat{\mathcal{V}}^{(\alpha)}\big\rangle_{h}\!-\!\big\langle\hat{\mathcal{V}}^{(\alpha)}\big\rangle_{h(\alpha)}\right).$ (15)
As $c^{(\alpha)}_{h}\!>\!c^{(\alpha)}_{l}$ for all $\alpha$ by definition, the first bracketed term is always positive. Hence, a sufficient condition for multi-parameter enhancement of the net work is that $\big\langle\hat{\mathcal{V}}^{(\alpha)}\big\rangle_{h}\!>\!\big\langle\hat{\mathcal{V}}^{(\alpha)}\big\rangle_{h(\alpha)}$.
Although the net work can be expressed as a sum of contributions associated with each control parameter, the corresponding expectation values are not independent. Each expectation value depends on all control parameters through the many-body state (i.e., the many-body density operator used for evaluating the expectation values). Consequently, varying one parameter modifies the expectation values entering the work contribution of another parameter. The enhancement therefore arises from this mutual dependence of thermodynamic observables on multiple control parameters, rather than from inherently quantum resources such as coherence. In this sense, the apparent deviation from simple additivity reflects the fact that the relevant degrees of freedom are not independent.
We note that the multi-parameter enhancement mechanism discussed here does not rely on uniquely quantum features. Rather, it originates from the dependence of thermodynamic observables on multiple control parameters, such that changing one parameter modifies the expectation values associated with another. This mechanism is therefore, in principle, applicable also to classical interacting systems. Quantum many-body systems, however, provide a natural setting where such cross-dependencies are strong and experimentally accessible.
III.1 Small quench sizes
When studying quantum Otto cycles, it is often useful to examine the case of small quench sizes, as such scenarios are commonly tractable to analytic methods. We therefore begin by defining the infinitesimal quench of the parameters $c^{(\alpha)}$ as $c^{(\alpha)}_{h}\!-\!c^{(\alpha)}_{l}\!\equiv\!dc^{(\alpha)}$. As noted previously, to have a multi-parameter enhancement to the net work, $\Delta W\!>\!0$, one must be able to demonstrate that the second bracketed term given in Eq. (15) is positive. For small quenches, we may re-express this difference as,
$\big\langle\hat{\mathcal{V}}^{(\alpha)}\big\rangle_{h}-\big\langle\hat{\mathcal{V}}^{(\alpha)}\big\rangle_{h(\alpha)}\simeq\sum_{\beta\neq\alpha}dc^{(\beta)}\frac{\partial\langle\hat{\mathcal{V}}^{(\alpha)}\big\rangle_{h}}{\partial c^{(\beta)}},$ (16)
where we have utilized the fact that the difference between these terms arises from the change in all parameters except $c^{(\alpha)}$. Substituting this expression into our formula for net work enhancement, given in Eq. (15), we find the approximate expression to second order,
$\Delta W\simeq\sum_{\alpha}dc^{(\alpha)}\sum_{\beta\neq\alpha}dc^{(\beta)}\frac{\partial\langle\hat{\mathcal{V}}^{(\alpha)}\rangle_{h}}{\partial c^{(\beta)}}.$ (17)
In the thermodynamic limit, the Hellmann-Feynman theorem allows for expression of the expectation value in a thermal equilibrium state as [52, 53, 54]
$\langle\hat{\mathcal{V}}^{(\alpha)}\rangle_{h}=\frac{\partial F_{h}}{\partial c^{(\alpha)}},$ (18)
where $F_{h}$ is the Helmholtz free energy in the high energy equilibrium state. We may therefore re-express the net work enhancement for small parameter quenches as
$\Delta W=2\sum_{\alpha<\beta}dc^{(\alpha)}dc^{(\beta)}\frac{\partial^{2}F_{h}}{\partial c^{(\alpha)}\partial c^{(\beta)}},$ (19)
where we have used the commutativity of partial derivatives to combine all pairs of mixed partial derivatives.
The second derivative of the free energy is directly related to the static susceptibility [46, 55]. As such, an interesting future direction would therefore be to extend the above to finite-time dynamics through linear response theory. In this paper though, we will utilize these small parameter quench results in Sec. VII for obtaining additional analytic insights for the TFIM, where the Helmholtz free energy at finite temperature, and hence its derivatives with respect to the Hamiltonian parameters, are know analytically.
IV The 1D Bose gas
@@
The Lieb-Liniger model of a 1D Bose gas is a paradigmatic quantum many-body model describing an ultracold atomic gas with repulsive contact pairwise interactions, being both experimentally realizable and having exact theoretical results [56, 57, 58, 7, 59, 60, 49, 58, 61, 62]. This model, in the uniform limit, consists first of an uncontrolled Hamiltonian operator, $\hat{H}_{0}$, corresponding to the kinetic energy,
@@
$c^{(2)}\hat{\mathcal{V}}^{(2)}=\frac{g}{2}\hat{\overline{G_{2}}}=\frac{g}{2}\!\int dx\hat{G}_{2}(x),$ (21)
where the two-body operator $\hat{\mathcal{V}}^{(2)}\equiv\hat{\overline{G_{2}}}=\int dx\hat{G}_{2}(x)$ corresponds to the integrated local (same-point) pair correlation $\hat{G}_{2}(x)\equiv\hat{G}_{2}(x,x)\!=\!\hat{\Psi}^{\dagger}(x)\hat{\Psi}^{\dagger}(x)\hat{\Psi}(x)\hat{\Psi}(x)$, which we note does not depend on $x$ in a uniform system due to the translational invariance. The interaction strength $c^{(2)}$ (with $c^{(2)}>0$ for repulsive interactions) is given by $c^{(2)}\!\equiv\!g/2\simeq\!\hbar\omega_{\perp}a_{s}$ away from confinement induced resonances [63] and is experimentally controllable via the tight transverse trapping frequency, $\omega_{\perp}$, or magnetic Feshbach resonance [64] tuning of the 3D $s$-wave scattering length $a_{s}$. We may utilize the normalized local (same-point) two-body correlation function,
@@
$c^{(1)}\hat{\mathcal{V}}^{(1)}\equiv\frac{1}{2}m\omega^{2}\int dx\,x^{2}\hat{\rho}(x),$ (23)
with $\hat{\mathcal{V}}^{(1)}\!\equiv\!\frac{1}{2}m\int dx\,x^{2}\hat{\rho}(x)$, where $c^{(1)}\!\equiv\!\omega^{2}$ corresponds to the harmonic trapping frequency squared and $\hat{\rho}(x)\!=\!\hat{\Psi}^{\dagger}(x)\hat{\Psi}(x)$ is the particle number density operator. We additionally define the atomic position variance, $\langle x^{2}\rangle\!\equiv\!\int dx\,x^{2}\langle\hat{\rho}(x)\rangle$, such that $\langle\hat{\mathcal{V}}^{(1)}\rangle\!=\!m\langle x^{2}\rangle/2$.
In the presence of an external trapping potential, one may again utilize the numerically exact TBA introduced above, along with the additional assumption of a local density approximation [54], to accurately model the non-uniform system at finite temperature. Through the external harmonic trap, the density profile gains a dependence on the position, $\rho\!=\!\rho(x)$. The gas may then be described in terms of the dimensionless interaction strength and temperature at the trap center, i.e. $\gamma_{0}\!=\!mg/\hbar^{2}\rho(0)$ and $\tau_{0}\!=\!2mk_{B}T/\hbar^{2}\rho(0)^{2}$, and the average total number of particles in the system, $N\!=\!\int dx\langle\hat{\rho}(x)\rangle$ [54]. The interaction-driven and separately the volumetric (i.e. harmonic frequency quench) quantum Otto cycles for this system have been previously investigated under both quasistatic and sudden quench protocols [20, 21, 27, 17, 16].
@@
IV.1 Two-parameter enhancement
Before turning to explicit calculation of the net work and efficiency, we provide a demonstration of the general multi-parameter enhancement to the net work for this two-parameter quantum Otto cycle. For clarity, we emphasize the dependence of the equilibrium states $(h)$ and $(l)$ on the interaction strength and harmonic trapping frequency as $\langle\cdot\rangle_{(g,\omega)}$. Using this notation, the formula for multi-parameter enhancement of the net work, $\Delta W$, given in Eq. (15), is expressed for the harmonically trapped 1D Bose gas as,